Solve by completing the square:
x^2 + 8x + 13 = 0

Answer by J2R2R(94)

Completing the square requires half the coefficient of x to be used (when the coefficient of x^2 is 1) so that we have a square on one side and since this isn’t 0 generally we have it equal to whatever we have to adjust the equation to which is 3 in this example.

(x + 4)^2 = x^2 + 8x + 16 so

x^2 + 8x + 13 = (x^2 + 8x + 16) - 3 = 0

(x + 4)^2 - 3 = 0

(x + 4)^2 = 3

Therefore

x + 4 = + or - square root of 3

x = - 4 + or - square root of 3